OpenStudy (anonymous):

solve the following equation u + 2 = 4 can you show all the steps to getting the answer

OpenStudy (anonymous):

what number would you add 2 to in order to get 4?

OpenStudy (anonymous):

well 2 but like I need to show my work and I know the answer but shouldn't it be more complicated

OpenStudy (austinl):

Nope. To get u by itself, just subtract 2 from each side. u=2

OpenStudy (anonymous):

can you guys help me with a couple of these please?

OpenStudy (anonymous):

-5 = s - 3?

OpenStudy (austinl):

Add 3 to each side. s=-5+3=?

OpenStudy (anonymous):

im not sure im confused I tried it and it got this wacko answer

OpenStudy (austinl):

-5 + 3 = ?

OpenStudy (anonymous):

-2

OpenStudy (austinl):

s=-2 :) I have to go now, hopefully someone else can help you!

OpenStudy (anonymous):

thank you for helping me as much as you did though :D

OpenStudy (anonymous):

5/13 t = -9

OpenStudy (wolfe8):

Since you multiply t with 5/13 to get -9, to just have t on the left side you will divide both sides with 5/13

OpenStudy (anonymous):

-117/5

OpenStudy (wolfe8):

Yup. If required or if you want to, you can write that as a mixed number or decimal. Usually mixed number.

OpenStudy (anonymous):

can you help me with two more?

OpenStudy (wolfe8):

Sure but can you make a new post for it? This one is getting long and the previous people would get notifications. It might bug them.

OpenStudy (anonymous):

sure thing ill link you